OpenStudy (anonymous):

what is an equation of the line that passes through the point (3, -8) and has the slope -2?

OpenStudy (anonymous):

equation of the line is given by (y-y1)=m(x-x1) , where 'm' is the slope and (x1,y1) is the point through which the line passes so we get... y+8=-2(x-3) on solving,we get 2x+y+2=0